Skip to content
Solved Assignments

Stat200 Introduction to Statistics

Fundamental concepts and methods in statistical analysis

Preview 5 of 14 pages

See the first 5 pages exactly as they appear in the document.

  1. Page 1
    Read page 1 as text

    Page 1 of 14 Stat200Introduction to Statistics Nam e______________________________ Answer Sheet Record your answers and work. Problem Number Solution 1 (25 pts) Answers: (a) True (b) True (c) False (d) True (e) False Work for (a), (b), (c), (d) and (e): 2 (5 pts) Answer: 70% Work: Checkout Time Frequency (in minutes) 1.0 - 1.9 2 2.0 - 2.9 12 3.0 - 3.9 2 4.0 - 4.9 4 T otal 20 percentage of the checkout times was less than 3 minutes = 14/20 = 0.7 = 70% 3 (5 pts) Answer: 2 nd class: 2.0 - 2.9 Work: Here, n = 20, Median = n/2th observation when n = even, Here, n = 20/2th obs = 10 th observation . Thus median is expected to lie in 2 nd class which is 2.0 – 2.9.

  2. Page 2
  3. Page 3
    Read page 3 as text

    Page 2 of 14 4 (5 pts) Answer: 2.85 Work: Checkout Time Frequency (in minutes) fi mid point (xi( fixi 1.0 - 1.9 2 1.45 2.9 2.0 - 2.9 12 2.45 29.4 3.0 - 3.9 2 3.45 6.9 4.0 - 4.9 4 4.45 17.8 T otal 20 57 Mean = ∑fixi / ∑fi = 57 / 20 = 2.85 5 (10 pts) Answer: 0.92 Work: Checkout Time Frequency (in minutes) F i mid point (xi( fixi fi*xi^2 1.0 - 1.9 2 1.45 2.9 4.205 2.0 - 2.9 12 2.45 29.4 72.03 3.0 - 3.9 2 3.45 6.9 23.805 4.0 - 4.9 4 4.45 17.8 79.21 20 57 179.25 Variance = (1/n)*{ ∑ (fi*xi^2) – n*x - bar^2 } = (1/20)*( 179.25 – 20*2.85^2) = (1/20)*(179.25 – 162.45) = (1/20)*(16.8) = 0.84 Standard deviation = sqrt(variance) = sqrt(0.84) = 0.9165 = 0.92 (approximately)

  4. Page 4
    Read page 4 as text

    Page 3 of 14 6 (10 pts) Ans wer: 13.77 Work: xi (xi - xbar)^2 2 256 15 9 15 9 18 0 40 484 total 90 758 Mean = = 90/5 =18 Sample Variance = = 758 / 4 =189.5 Standard deviation = sqrt(variance) = sqrt(189.5) = 13.7659 = 13.77 (approximately) 7 (5 pts) Answer: 0.765 Work: Coefficient of variation = = =0.765 8 (5 pts) Answer: no Work:

  5. Page 5
    Read page 5 as text

    Page 4 of 14 9 (10 pts) Answer: 1/221 Work: Probability (1 st card is ace and 2 nd card is also an ace) = ? Probability of drawing 1 st card as ace = 4/52 Probability of drawing 2 nd card as ace without replacement = 3/51 Thus required probability = (4/52)*(3/51) = (1/13)*(1/17) = 1/221 10 (10 pts) Answer: 1/169 Work: Probability (1 st card is ace and 2 nd card is also an ace) = ? Probability of drawing 1 st card as ace = 4/52 Probability of drawing 2 nd card as ace with replacement = 4/52 Thus required probability = (4/52)*(4/52) = (1/13)*(1/13) = 1/169 11 (5 pts) Answer: yes, events are independent in case of with replacement because P(A) * P(B) = P(A and B) Work: two events are said to be independent if P(AandB) = P(A) * P(B) In case of with replacement, LetA denote the event of drawing 1 st ace and event B denote the event of drawing 2 nd ace. Now, P(A) = Probability of drawing 1 st card as ace = 4/52 = 1/13 P(B) = Probability of drawing 2 nd card as ace (with replacement) = 4/52 = 1/13 P(A) * P(B) = 1/13 * 1/13 = 1/169 P(A and B) = P(drawing 1 st card as ace and 2 nd card as ace without replacement) = (4/52 * 4/52) = 1/169 Since, P(A and B) = P(A)*P(B), thus events are independent 12 (10 pts) Answer: 0.25 Work: Let event A denote STAT200 And event B denote PHY300 Here total number of juniors in college is 1000

9 more pages in the full document

Unlock it once and it stays in your library, ready to chat with or turn into flashcards.

Unlock the full document
See all

Study this document with CramX

Once it is in your library, every tool can work from it.